OpenStudy (cutiecomittee123):

the value pi/2 is a solution to 4cos^2 (4x)-3=0 true or false?

OpenStudy (campbell_st):

ok... so solve it and you get 4cos^2(4x) = 3 divde by 4 \[\cos^2(4x) = \frac{3}{4}\] next take the square roots \[\cos(4x) = \pm \frac{\sqrt{3}}{2}\] now you should be able o solve from here as you have an exact value. when you get the 2 angles, divide them both by 4... to get x.

OpenStudy (campbell_st):

hope it helps

OpenStudy (campbell_st):

lol.... opps.... just substitute pi/2 into the equation. you don't need to solve it....

OpenStudy (cutiecomittee123):

i already chose false on it as a wild guess to keep going. yikes.

OpenStudy (campbell_st):

so when substituting you get \[4 \cos^2(4 \times \frac{\pi}{2}) - 3 = 0\] which is \[4\cos^2(2\pi) - 3 = 0\] so find cos(2pi) and work from there.

OpenStudy (campbell_st):

so I'd say false is the correct choice.

OpenStudy (cutiecomittee123):

thank goodness. lol Well I am stuggling a little with another, can you help again lol?